How many digits are used for printing the page numbers of a book that has 163 pages?
Answer by ikleyn(52776)

.
9 pages have one-digit number (1, 2, 3, . . . , 9).  1*9  = 9 digits.

90 pages have two-digit numbers (10, 11, 12, . . . , 99).  2*90 = 180 digits.


64 remaining pages have three-digit numbers (100, 101, . . . , 164).  3*64 = 192 digits.

Your answer is 9 + 99 + 192.